# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

### v2

@@ III-D Local Keypoint Feature Aggregator
As illustrated in Fig. 1(c), we first design a coordinate encoding function that projects the keypoint coordinates into a high-dimensional feature space, producing positional embeddings $\mathcal{E}_{kpt}\in\mathbb{R}^{N_{kpt}\times d}$. These embeddings are then added to the initial keypoint features $\mathcal{F}_{kpt}$ to enhance spatial awareness:
$\mathcal{E}_{kpt}=MLP(\mathcal{P}_{kpt}),\quad\mathcal{F}_{kpt}=\mathcal{F}_{kpt}+\mathcal{E}_{kpt}.$ (5)
@@
$\mathcal{F}_{K}^{\prime}=Norm(Concat(\mathcal{F}_{K},\mathcal{P}_{K})).$ (6)
@@
$\mathcal{F}_{K}^{\prime\prime}=Concat(\mathcal{F}_{K}^{\prime},\mathcal{F}_{kpt}),\quad\hat{\mathcal{F}}_{K}=\mathbf{\alpha}\odot\mathcal{F}_{K}^{\prime\prime}+\mathbf{\beta},$ (7)
@@
$\mathcal{F}_{local}=MLP(MaxPool(\mathbf{\hat{\mathcal{F}}}_{K})).$ (8)
@@ III-E Global Keypoint Feature Aggregator
To complement the geometrically distinct local features captured by the LKFA with geometrically consistent global features across different object instances, we introduce a Global Keypoint Feature Aggregator (GKFA) based on SSMs. This design facilitates robust modeling of global shape structure and improves cross-instance generalization at the category level. To clarify the theoretical underpinnings of our GKFA module, we begin by reviewing the core formulation of SSMs, followed by an introduction to the Mamba architecture built upon them, which serves as the foundation of our global feature modeling.
III-E1 Preliminaries
SSMs [31] were originally introduced in control theory for modeling linear time-invariant (LTI) dynamical systems. The core principle is to map the input signal $x(t)$ to an output signal $y(t)$ through a latent state $h(t)$, whose evolution is governed by both the current state and the input. This process is described by the following system of ordinary differential equations:
$h^{\prime}(t)=Ah(t)+Bx(t),\quad y(t)=Ch(t),$ (9)
where $A$ is the state transition matrix, and $B$ and $C$ are the input and output projection matrices, respectively. Here, $t$ denotes continuous time, and the system captures how the internal dynamics evolve smoothly over time.
To apply SSMs to discrete data such as point cloud sequences or tokens in deep learning, the continuous-time formulation must be discretized. Mamba [32] achieves this via the zero-order hold method, which introduces a time step ($\Delta$) to approximate the continuous dynamics. This leads to a transformation of the continuous matrices $A$ and $B$ into their discrete counterparts $\overline{A}$ and $\overline{B}$, as follows:
$\overline{A}=e^{\Delta A},\quad\overline{B}=(\Delta A)^{-1}(e^{\Delta A}-I)\Delta B.$ (10)
Here, $\Delta$ represents the temporal resolution of the sampled sequence (i.e., the interval between adjacent observations in discrete time). With these discretized parameters, the continuous-time system in Eq. (9) can be reformulated into a recurrent state-space model defined over discrete time steps:
$h_{t}=\overline{A}h_{t-1}+\overline{B}x_{t},\quad y_{t}=Ch_{t},$ (11)
where $t$ denotes the discrete time step. This formulation enables efficient sequence-to-sequence mapping from inputs $x_{t}$ to outputs $y_{t}$, but its inherently sequential nature poses challenges for efficient training.
To support parallelized training, the recurrence can be equivalently expressed in convolutional form:
$\overline{K}=(C\overline{B},C\overline{AB},\ldots,C\overline{A}^{L-1}\overline{B}),\quad y=x\ast\overline{K},$ (12)
where $L$ denotes the length of the input sequence $x$, and $\overline{K}$ represents the global convolution kernel composed of weighted combinations of the discretized system parameters. The operator * denotes the one-dimensional convolution, which applies the kernel $\overline{K}$ over the input sequence $x$ to generate the output $y$. This convolutional formulation allows the model to compute the entire output sequence in parallel, significantly improving training efficiency over the recursive formulation.
III-E2 GKFA Module
As illustrated in Fig. 1(d), to capture global semantic consistency and long-range dependencies among keypoints, we integrate a bidirectional variant of Mamba [35] into the GKFA module. This design enables bidirectional information flow across keypoints, facilitating the learning of globally consistent geometric structures. Traditionally, bidirectional modeling is achieved through a strategy we refer to as Point Sequence Flipping (PSF), where the input point sequence is reversed based on spatial indices. However, such input-order-based reversal is not well suited for point cloud data, as the unordered and non-causal nature of point sets may disrupt spatial and semantic coherence, especially when geometrically unrelated points are juxtaposed in the reversed sequence.
To address this limitation, we propose a Feature Sequence Flipping (FSF) strategy that operates along the feature dimension rather than spatial order. Unlike PSF, FSF preserves the original spatial arrangement of keypoints, while enabling bidirectional modeling through the feature space. As illustrated in Fig. 3, we construct a forward feature sequence by directly using the keypoint features $\mathcal{F}_{local}$, and generate a backward feature sequence by reversing the feature channels of $\mathcal{F}_{local}$. These two sequences are then independently fed into two parallel Mamba blocks. The outputs are aggregated via element-wise summation, followed by a linear projection and residual connection, to produce the final globally aggregated keypoint features $\mathcal{F}_{global}\in\mathbb{R}^{N_{kpt}\times d}$:
$\mathcal{F}_{global}=SSM(\mathcal{F}_{local},FSF(\mathcal{F}_{local}))+\mathcal{F}_{local}.$ (13)
Figure 3: Comparison between traditional Point Sequence Flipping (PSF) and the proposed Feature Sequence Flipping (FSF).
@@ III-E2 GKFA Module
To ensure that the aggregated keypoints retain meaningful and expressive geometric structure, we introduce a reconstruction-based supervision mechanism. Specifically, we employ an MLP layer to regress dense positional offsets from the global keypoint features. Given $\mathcal{F}_{global}\in\mathbb{R}^{N_{kpt}\times d}$, we repeat each feature vector $m=\lfloor N_{rec}/N_{kpt}\rfloor$ times to construct an expanded representation $\mathcal{F}_{global}^{\prime}\in\mathbb{R}^{N_{rec}\times d}$, where $N_{rec}$ is the predefined number of reconstruction points (set to 960 in our experiments). The MLP predicts a set of 3D offsets $\mathcal{O}_{\mathcal{P}}\in\mathbb{R}^{N_{rec}\times 3}$:
$\mathcal{O}_{\mathcal{P}}=MLP(\mathcal{F}_{global}^{\prime}).$ (14)
@@
Similarly, the keypoint coordinates $\mathcal{P}_{kpt}\in\mathbb{R}^{N_{kpt}\times 3}$ are repeated $m$ times to obtain $\mathcal{P}_{kpt}^{\prime}\in\mathbb{R}^{N_{rec}\times 3}$. The reconstructed point cloud $\mathcal{P}_{rec}$ is then computed as:
$\mathcal{P}_{rec}=\mathcal{P}_{kpt}^{\prime}+\mathcal{O}_{\mathcal{P}}.$ (15)
@@
This reconstruction supervision encourages each keypoint to act as a generator of local geometry, effectively bridging the gap between sparse semantic abstraction and dense geometric representation. In our ablation studies, we compare different settings for the number of keypoints and find that using 96 (with a repetition factor of 10) achieves the best performance under a fixed reconstruction resolution of 960 points.
To further strengthen the representational capacity of keypoints, we stack multiple LKFA and GKFA modules in a cascaded architecture. Each block refines keypoint features by alternately aggregating fine-grained local geometry and long-range global context. This design allows hierarchical learning across different scales, improving the network’s ability to model complex object structures. As demonstrated by our ablation studies (Section IV-C), using $S$ = 12 stacked blocks yields the best performance.
@@ III-F Pose and Size Estimator
After obtaining semantically consistent and geometrically informative keypoints along with their associated features, we develop a Pose and Size Estimator to directly infer the object’s 6D pose and size. Specifically, we first map the predicted keypoint coordinates to the NOCS space using the ground-truth rotation $\textit{R}_{gt}$, translation $\textit{t}_{gt}$, and scale $\textit{s}_{gt}$, thereby obtaining the ground-truth keypoint coordinates in NOCS space $\mathcal{P}_{kpt}^{gt}\in\mathcal{R}^{N_{kpt}\times 3}$. Subsequently, we predict the NOCS coordinates $\mathcal{P}_{kpt}^{nocs}\in\mathcal{R}^{N_{kpt}\times 3}$ via a self-attention mechanism followed by an MLP layer:
$\mathcal{P}_{kpt}^{gt}=\left\{\frac{1}{\|\textit{s}_{gt}\|_{2}}(x_{i}-\textit{t}_{gt})\textit{R}_{gt}\mid x_{i}\in\mathcal{P}_{kpt}\right\},$ (17)
$\mathcal{P}_{kpt}^{nocs}=MLP(Self\text{-}Attn(\mathcal{F}_{kpt})).$ (18)
To supervise the mapping from keypoints to NOCS, we design a mapping loss $\mathcal{L}_{map}$, employing the smooth $\mathcal{L}_{1}$ loss [14] to constrain the distance between the predicted NOCS coordinates and the ground-truth NOCS coordinates.
By leveraging keypoint-level correspondences, we employ three parallel MLPs to predict the rotation $R_{pre}$, translation $t_{pre}$ and size $s_{pre}$, respectively.
$\mathcal{F}_{pose}=Concat(\mathcal{P}_{kpt},\mathcal{P}_{kpt}^{nocs},\mathcal{F}_{kpt}),$ (19)
$R_{pre},t_{pre},s_{pre}=MLP(\mathcal{F}_{pose}),MLP(\mathcal{F}_{pose}),MLP(\mathcal{F}_{pose}),$ (20)
where $R_{pre}$ and $t_{pre}$ represent the predicted 6D object pose, and $s_{pre}$ represents the predicted object size. We simply use the $\mathcal{L}_{2}$ norm loss function to supervise $R_{pre}$, $t_{pre}$ and $s_{pre}$:
@@ IV-A2 Evaluation Metrics
TABLE I: Quantitative comparisons With state-of-the-art methods on the REAL275 dataset[10]. The best results are in bold.
  Method Prior REAL275 ($\uparrow$)
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$
SPD [ECCV’20][11] 77.3 53.2 19.3 21.4 43.2 54.1
SGPA [ICCV’21][15] 80.1 61.9 35.9 39.6 61.3 70.7
GPV-Pose [CVPR’22][41] 83.0 64.4 32.0 42.9 55.0 73.3
DPDN [ECCV’22][16] 83.4 76.0 46.0 50.7 70.4 78.4
GPT-COPE [TCSVT’23][18] 82.0 70.4 45.9 53.8 63.1 77.7
HS-Pose [CVPR’23][42] 82.1 74.7 46.5 55.2 68.6 82.7
IST-Net [ICCV’23][13] 82.5 76.6 47.5 53.4 72.1 80.5
Query6DoF [ICCV’23][20] 82.5 76.1 49.0 58.9 68.7 83.0
VI-Net [ICCV’23][22] - - 50.0 57.6 70.8 82.1
CatFormer [AAAI’24][19] 83.1 73.8 47.7 53.7 69.0 79.5
CLIPose [TCSVT’24][12] - - 48.7 58.3 70.4 85.2
TG-Pose [TMM’24][21] 82.6 76.2 49.8 59.0 71.7 86.6
SecondPose [CVPR’24][23] - - 56.2 63.6 74.7 86.0
AG-Pose [CVPR’24][14] 83.8 77.6 56.2 62.3 73.4 81.2
AG-Pose (DINOv2) [CVPR’24][14] 84.1 80.1 57.0 64.6 75.1 84.7
SpherePose [ICLR’25][29] 84.0 79.0 58.2 67.4 76.2 88.2
Ours 84.0 80.6 63.2 68.3 80.9 88.0
TABLE II: Quantitative comparisons with state-of-the-art methods on the CAMERA25 dataset[10]. The best results are in bold.
  Method Prior CAMERA25 ($\uparrow$)
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$
SPD [ECCV’20] [11] 93.2 83.1 54.3 59.0 73.3 81.5
SGPA [ICCV’21] [15] 93.2 88.1 70.7 74.5 82.7 88.4
GPV-Pose [CVPR’22] [41] 92.9 86.6 67.4 76.2 - 87.4
GPT-COPE [TCSVT’23] [18] 92.5 86.9 70.4 76.5 81.3 88.7
HS-Pose [CVPR’23] [42] 93.3 89.4 73.3 80.5 80.4 89.4
Query6DoF [ICCV’23] [20] 91.9 88.1 78.0 83.1 83.9 90.0
VI-Net [ICCV’23] [22] - - 74.1 81.4 79.3 87.3
TG-Pose [TMM’24] [21] 93.4 89.7 74.5 81.9 81.1 90.4
CatFormer [AAAI’24] [19] 93.5 89.9 74.9 79.8 85.3 90.2
CLIPose [TCSVT’24] [12] - - 74.8 82.2 82.0 91.2
AG-Pose [CVPR’24] [14] 94.1 91.7 77.1 82.0 85.5 91.6
AG-Pose (DINOv2) [CVPR’24] [14] 94.2 92.5 79.5 83.7 87.1 92.6
SpherePose [ICLR’25] [29] 94.8 92.4 78.3 84.3 84.8 92.3
Ours 94.4 92.9 81.0 85.0 88.1 93.4
Figure 4: Qualitative comparisons between INKL-Pose with AG-Pose[14] on the REAL275 dataset[10], where green and red represent the predicted and ground-truth bounding boxes, respectively. Our INKL-Pose demonstrates superior precision in pose estimation.
@@ IV-A3 Implementation Details
Regarding optimization, the weights of the loss terms are set as {$w_{sep},w_{surf},w_{sim},w_{map},w_{pose}$} = {10.0, 10.0, 15.0, 2.0, 0.3}. The model is trained using the Adam optimizer [44] with a cyclical learning rate schedule [45], where the learning rate varies between $2\times 10^{-5}$ and $5\times 10^{-4}$. All experiments are conducted on a single NVIDIA RTX 4090D GPU with a batch size of 64. For CAMERA25 and REAL275, we train the model using both synthetic and real-world data with a synthetic-to-real ratio of 3:1. For HouseCat6D, the model is trained exclusively on its corresponding dataset.

### v3

@@ III-D Local Keypoint Feature Aggregator
As illustrated in Fig. 3(a), we first utilize a coordinate encoding function, implemented as an MLP layer, to project the keypoint coordinates into a high-dimensional feature space, producing positional embeddings $\mathcal{E}_{kpt}\in\mathbb{R}^{N_{kpt}\times d}$. These embeddings are then added to the initial keypoint features $\mathcal{F}_{kpt}$ to enhance spatial awareness:
$\mathcal{E}_{kpt}=\text{MLP}(\mathcal{P}_{kpt}),\quad\mathcal{F}_{kpt}=\mathcal{F}_{kpt}+\mathcal{E}_{kpt}.$ (5)
@@
$\mathcal{F}_{K}^{\prime}=\text{Norm}(\text{Concat}(\mathcal{F}_{K},\mathcal{P}_{K})).$ (6)
@@
$\mathcal{F}_{K}^{\prime\prime}=\text{Concat}(\mathcal{F}_{K}^{\prime},\mathcal{F}_{kpt}),\quad\hat{\mathcal{F}}_{K}=\mathbf{\alpha}\odot\mathcal{F}_{K}^{\prime\prime}+\mathbf{\beta},$ (7)
@@
$\mathcal{F}_{local}=\text{MLP}(\text{MaxPool}(\mathbf{\hat{\mathcal{F}}}_{K})).$ (8)
@@ III-E Global Keypoint Feature Aggregator
While the Local Keypoint Feature Aggregator (LKFA) effectively captures fine-grained geometric details, it lacks the capacity to model global structural context and long-range dependencies among keypoints. To address this limitation, we introduce a Global Keypoint Feature Aggregator (GKFA), which enhances semantic consistency by leveraging SSMs.
As illustrated in Fig. 3(b), GKFA employs a bidirectional Mamba-based architecture [26] to efficiently model long-range dependencies and capture object-level structural information. Compared to traditional Transformers, which suffer from quadratic complexity with respect to sequence length, Mamba operates in linear time, making it a more scalable and efficient choice for processing keypoint sequences in point clouds.
To fully exploit Mamba’s bidirectional modeling capabilities, it is necessary to construct both forward and backward feature sequences. A common practice in prior works is to reverse the input sequence order, a technique we refer to as Point Sequence Flipping (PSF). However, point cloud data lack an inherent spatial order, and reversing the index sequence may place geometrically unrelated keypoints next to each other, thereby disrupting the continuity of feature learning. To address this issue, we propose a simple yet effective strategy termed Feature Sequence Flipping (FSF). Rather than reordering keypoints, FSF reverses the feature channels within each keypoint vector while preserving the original spatial arrangement of the point cloud. Formally, given a keypoint feature vector $\mathcal{F}_{local}(i)\in\mathbb{R}^{d}$ expressed as $[f_{1}(i),f_{2}(i),\ldots,f_{d}(i)]$, we define FSF as:
$\text{FSF}(\mathcal{F}_{local}(i))=[f_{d}(i),f_{d-1}(i),\ldots,f_{1}(i)].$ (9)
Since different feature channels encode complementary semantics such as geometry or part cues, this reversal introduces a novel perspective on feature dependencies while maintaining the geometric integrity of the point cloud.
In the GKFA, both the original (forward) and FSF-processed (backward) feature sequences are fed into two parallel Mamba blocks. Each block independently captures long-range dependencies within its respective sequence. The outputs from the two Mamba blocks are then aggregated via element-wise summation, forming the final global keypoint features:
$\mathcal{F}_{global}=\text{SSM}(\mathcal{F}_{local},\text{FSF}(\mathcal{F}_{local}))+\mathcal{F}_{local}.$ (10)
This design allows the model to incorporate both the original and flipped feature perspectives, promoting more comprehensive global context learning while maintaining computational efficiency and stability.
@@
To ensure that the aggregated keypoints retain meaningful and expressive geometric structure, we utilize the reconstruction-based supervision mechanism [20]. Specifically, we employ an MLP layer to regress dense positional offsets from the global keypoint features. Given $\mathcal{F}_{global}\in\mathbb{R}^{N_{kpt}\times d}$, we repeat each feature vector $m=\lfloor N_{rec}/N_{kpt}\rfloor$ times to construct an expanded representation $\mathcal{F}_{global}^{\prime}\in\mathbb{R}^{N_{rec}\times d}$, where $N_{rec}$ is the predefined number of reconstruction points (set to 960 in our experiments). The MLP predicts a set of 3D offsets $\mathcal{O}_{\mathcal{P}}\in\mathbb{R}^{N_{rec}\times 3}$:
$\mathcal{O}_{\mathcal{P}}=\text{MLP}(\mathcal{F}_{global}^{\prime}).$ (11)
@@
Similarly, the keypoint coordinates $\mathcal{P}_{kpt}\in\mathbb{R}^{N_{kpt}\times 3}$ are repeated $m$ times to obtain $\mathcal{P}_{kpt}^{\prime}\in\mathbb{R}^{N_{rec}\times 3}$. The reconstructed point cloud $\mathcal{P}_{rec}$ is then computed as: $\mathcal{P}_{rec}=\mathcal{P}_{kpt}^{\prime}+\mathcal{O}_{\mathcal{P}}.$
@@
To further improve representational capacity, we stack multiple LKFA and GKFA modules, refining keypoint features hierarchically. Our ablation studies (Section IV-C) show that stacking $S=12$ blocks yields the best overall performance.
@@ III-F Pose and Size Estimator
After obtaining semantically consistent and geometrically informative keypoints and their associated features, we adopt a standard Pose and Size Estimator, as in previous work [14, 20], to infer the object’s 6D pose and size. Specifically, we first transform the predicted keypoint coordinates into the NOCS space using the ground-truth rotation $\textit{R}_{gt}$, translation $\textit{t}_{gt}$, and scale $\textit{s}_{gt}$:
$\mathcal{P}_{kpt}^{gt}=\left\{\frac{1}{\|\textit{s}_{gt}\|_{2}}(x_{i}-\textit{t}_{gt})\textit{R}_{gt}\mid x_{i}\in\mathcal{P}_{kpt}\right\}.$ (13)
We then predict the NOCS keypoints from the extracted features via a self-attention mechanism followed by an MLP layer:
$\mathcal{P}_{kpt}^{nocs}=\text{MLP}(\text{Self-Attn}(\mathcal{F}_{kpt})).$ (14)
To supervise this mapping, we adopt a smooth $\mathcal{L}_{1}$ loss:
$\mathcal{L}_{map}=\frac{1}{N_{kpt}}\sum_{i=1}^{N_{kpt}}\text{Smooth-}L_{1}\left(\mathcal{P}_{kpt}^{nocs}(i)-\mathcal{P}_{kpt}^{gt}(i)\right),$ (15)
[…]
Finally, we regress the object rotation $R_{pre}$, translation $t_{pre}$, and size $s_{pre}$ using parallel MLPs based on the concatenated keypoint information: $\mathcal{F}_{pose}=\text{Concat}(\mathcal{P}_{kpt},\mathcal{P}_{kpt}^{nocs},\mathcal{F}_{kpt})$. We apply an $\mathcal{L}_{2}$ loss to supervise these predictions:
@@ IV-A3 Implementation Details
Regarding optimization, the weights of the loss terms are set as {$w_{sep},w_{surf},w_{sim},w_{map},w_{pose}\}=\{10.0,10.0,15.0,2.0,0.3\}$. The model is trained using the Adam optimizer [33] with a cyclical learning rate schedule [34], where the learning rate varies between $2\times 10^{-5}$ and $5\times 10^{-4}$. All experiments are conducted on a single NVIDIA RTX 4090D GPU with a batch size of 64. For CAMERA25 and REAL275, we train the model using both synthetic and real-world data with a synthetic-to-real ratio of 3:1. For HouseCat6D, the model is trained exclusively on its corresponding dataset.
Method Prior REAL275 ($\uparrow$) #Params
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$ (M) $\downarrow$
SPD [ECCV’20][11] 77.3 53.2 19.3 21.4 43.2 54.1 18.3
SGPA [ICCV’21][13] 80.1 61.9 35.9 39.6 61.3 70.7 23.3
GPV-Pose [CVPR’22][35] 83.0 64.4 32.0 42.9 55.0 73.3 8.6
DPDN [ECCV’22][14] 83.4 76.0 46.0 50.7 70.4 78.4 24.6
IST-Net [ICCV’23][17] 82.5 76.6 47.5 53.4 72.1 80.5 26.8
Query6DoF [ICCV’23][15] 82.5 76.1 49.0 58.9 68.7 83.0 19.7
VI-Net [ICCV’23][18] - - 50.0 57.6 70.8 82.1 28.9
CatFormer [AAAI’24][12] 83.1 73.8 47.7 53.7 69.0 79.5 -
CLIPose [TCSVT’24][16] - - 48.7 58.3 70.4 85.2 -
SecondPose [CVPR’24][19] - - 56.2 63.6 74.7 86.0 38.1
AG-Pose [CVPR’24][20] 83.8 77.6 56.2 62.3 73.4 81.2 28.9
AG-Pose (DINOv2) [CVPR’24][20] 84.1 80.1 57.0 64.6 75.1 84.7 11.0
SpherePose [ICLR’25][21] 84.0 79.0 58.2 67.4 76.2 88.2 -
KeyPose [AAAI’25][36] 84.2 80.0 57.7 66.0 78.8 88.0 -
GCE-Pose [CVPR’25][37] 84.1 79.8 57.0 65.1 75.6 86.3 -
Ours 84.0 80.6 63.2 68.3 80.9 88.0 16.7
TABLE I: Quantitative comparisons with state-of-the-art methods on the REAL275 dataset[10]. The best results are in bold. The symbol indicates that only the trainable parameter count is reported. The symbol -- indicates that the data is not available or not reported.
Method Prior CAMERA25 ($\uparrow$) #Params
$\mathrm{IoU_{50}}$ $\mathrm{IoU_{75}}$ $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$ (M) $\downarrow$
SPD [ECCV’20] [11] 93.2 83.1 54.3 59.0 73.3 81.5 18.3
SGPA [ICCV’21] [13] 93.2 88.1 70.7 74.5 82.7 88.4 23.3
GPV-Pose [CVPR’22] [35] 92.9 86.6 67.4 76.2 - 87.4 8.6
Query6DoF [ICCV’23] [15] 91.9 88.1 78.0 83.1 83.9 90.0 19.7
VI-Net [ICCV’23] [18] - - 74.1 81.4 79.3 87.3 28.9
CatFormer [AAAI’24] [12] 93.5 89.9 74.9 79.8 85.3 90.2 -
CLIPose [TCSVT’24] [16] - - 74.8 82.2 82.0 91.2 -
AG-Pose [CVPR’24] [20] 94.1 91.7 77.1 82.0 85.5 91.6 28.9
AG-Pose (DINOv2) [CVPR’24] [20] 94.2 92.5 79.5 83.7 87.1 92.6 11.0
SpherePose [ICLR’25] [21] 94.8 92.4 78.3 84.3 84.8 92.3 -
KeyPose [AAAI’25][36] 94.4 92.6 79.8 83.6 87.1 92.3 -
Ours 94.4 92.9 81.0 85.0 88.1 93.4 16.7
TABLE II: Quantitative comparisons with state-of-the-art methods on the CAMERA25 dataset [10]. The best results are in bold. The symbol indicates that only the trainable parameter count is reported. The symbol -- indicates that the data is not available or not reported.
Figure 4: Qualitative comparisons between INKL-Pose with AG-Pose [20] on the REAL275 dataset, where green and red represent the predicted and ground-truth bounding boxes, respectively. Our INKL-Pose demonstrates superior precision in pose estimation.